# Ramification filtration via deformations, II

### v3

@@ 2.3. Lie algebras ${\mathcal{L}}^{\,{\dagger}}$ and $\bar{\mathcal{L}}^{\,{\dagger}}$
Let $\iota=p^{m}(qp^{M-1}\alpha-(q-1)\beta)\in\mathfrak{A}^{0}$ be given in the standard notation from Sect. 2.2.
@@
The elements of $\{t^{-\iota}\ |\ \iota\in\mathfrak{A}^{+}(p)\}$ behave “well” modulo $(\sigma-\mathrm{id})\mathcal{K}$, i.e. the natural map $\sum\limits_{\iota\in\mathfrak{A}^{+}(p)}kt^{-\iota}\longrightarrow\mathcal{K}/(\sigma-\mathrm{id})\mathcal{K}$ is injective. This is implied by the following proposition.
@@
Indeed, $\iota p^{-m}=qp^{M-1}\alpha-(q-1)\beta\in p^{M}\mathbb{N}$ implies (use that $qp^{M-1}\alpha\in p^{M}\mathbb{N}$ because $m<N^{*}$) that $\beta=p^{-m}(b_{1}p^{m_{1}}+b_{2}p^{m_{2}}+\dots+b_{u}p^{m_{u}})\in p^{M}\mathbb{N}$ with all $m_{i}\leqslant m+M-1$. But this number is $\leqslant p^{M-1}\mathrm{ch}\,\iota<p^{M}$. The contradiction. Then Prop. 2.1a) implies that
@@
Finally, if $\iota\in\mathfrak{A}^{0}_{0}$ then $\iota p^{-v_{p}(\iota)}=a<(p-1)v_{0}$ implies that $a<(v_{0}-\delta_{0})(p-1)$ by the choice of $\delta_{0}$, cf. Sect. 2.1. On the other hand, for all $\iota\in\mathfrak{A}^{+}(p)$ with $\mathrm{ch}(\iota)\geqslant 1$, the values $\iota p^{-v_{p}(\iota)}$ are different (use that $\mathrm{gcd}(p^{m}\alpha,p^{m}\beta/r^{*})\not\equiv 0\,\mathrm{mod}\,p$) and are bigger than $(v_{0}-\delta_{0})(p-1)$.
@@
Let ^$\widehat{\mathcal{L}}_{W_{M}(k)}^{\,{\dagger}}$ be the Lie algebra over $W_{M}(k)$ with the set of free generators
@@
Use notation $\sigma$ for the $\sigma$-linear automorphism of ^$\widehat{\mathcal{L}}^{\,{\dagger}}_{W_{M}(k)}$ such that $\sigma:D^{\,{\dagger}}_{\iota n}\mapsto D^{\,{\dagger}}_{\iota,n+1}$ and $\sigma:D^{{\dagger}}_{0}\mapsto D^{{\dagger}}_{0}$. Introduce the Lie $\mathbb{Z}/p^{M}$-algebras ^:=^$\widehat{\mathcal{L}}^{\,{\dagger}}:=\widehat{\mathcal{L}}^{\,{\dagger}}_{W_{M}(k)}|_{\sigma=\mathrm{id}}$ and =^$\mathcal{L}^{\,{\dagger}}=\widehat{\mathcal{L}}^{\,{\dagger}}/C_{p}(\widehat{\mathcal{L}}^{\,{\dagger}})$. Note that ^$\widehat{\mathcal{L}}^{\,{\dagger}}\otimes W_{M}(k)=\widehat{\mathcal{L}}^{\,{\dagger}}_{W_{M}(k)}$, and this matches the agreement about extensions of scalars from the Introduction.
Denote by $\{\mathcal{L}^{\,{\dagger}}(w)\}_{w\geqslant 1}$ the minimal central filtration of $\mathcal{L}^{\,{\dagger}}$ such that $p^{m}D^{\,{\dagger}}_{\iota n}\in\mathcal{L}^{\,{\dagger}}(w)_{W_{M}(k)}$ if $\iota\geqslant p^{M-1}b^{*}U(w-1,m)$.
This means that $\mathcal{L}^{\,{\dagger}}(w)_{W_{M}(k)}$ is an ideal in $\mathcal{L}^{\,{\dagger}}_{W_{M}(k)}$ generated as $W_{M}(k)$-module by all $p^{m}[\dots[D^{\,{\dagger}}_{\iota_{1}n_{1}},D^{\,{\dagger}}_{\iota_{2}n_{2}}],\dots,D^{\,{\dagger}}_{\iota_{r}n_{r}}]$ such that for some integers $m_{1},\dots,m_{r},s_{1},\dots,s_{r}\geqslant 0$ we have:
a) $\iota_{j}\geqslant p^{M-1}b^{*}U(s_{j},m_{j})$, where $j=1,\ldots,r$;
b) $r+\sum_{1\leqslant j\leqslant r}s_{j}\geqslant w$ and $\sum_{1\leqslant j\leqslant r}m_{j}\leqslant m$.
Note that for any $w\geqslant 1$, $C_{w}(\mathcal{L}^{\,{\dagger}})\subset\mathcal{L}^{\,{\dagger}}(w)$.
For $m\in\mathbb{Z}_{\geqslant 0}$, let $\mathcal{L}^{\,{\dagger}}[m]_{W_{M}(k)}$ be the $W_{M}(k)$-submodule of $\mathcal{L}^{{\dagger}}_{W_{M}(k)}$ generated by all monomials $[D^{\,{\dagger}}_{\bar{\iota}\bar{n}}]:=[\ldots[D^{\,{\dagger}}_{\iota_{1}n_{1}},\ldots],D^{\,{\dagger}}_{\iota_{r}n_{r}}]$ such that $\mathrm{ch}(\iota_{1})+\dots+\mathrm{ch}(\iota_{r})=m$. Denote by $\mathcal{L}^{{\dagger}}[m]$ the $\mathbb{Z}_{p}$-submodule of $\mathcal{L}^{{\dagger}}$ such that $\mathcal{L}^{{\dagger}}[m]\otimes W_{M}(k)=\mathcal{L}^{{\dagger}}[m]_{W_{M}(k)}$.
@@
For $\bar{\iota}=(\iota_{1},\dots,\iota_{r})$ set $r(\bar{\iota})=r$, $\mathrm{ch}(\bar{\iota})=\mathrm{ch}(\iota_{1})+\dots+\mathrm{ch}(\iota_{r})$ and $|\bar{\iota}|=\iota_{1}+\dots+\iota_{r}$.
As a result, we have the grading $\mathcal{L}^{\,{\dagger}}=\oplus_{m\geqslant 0}\mathcal{L}^{\,{\dagger}}[m]$ in the category of Lie algebras. For $w\geqslant 1$, set $\mathcal{L}^{\,{\dagger}}(w)[m]=\mathcal{L}^{\,{\dagger}}(w)\cap\mathcal{L}^{\,{\dagger}}[m]$. Then $\mathcal{L}^{\,{\dagger}}(w)=\oplus_{m\geqslant 0}\mathcal{L}^{\,{\dagger}}(w)[m]$.
@@
The algebra $\mathcal{L}^{{\dagger}}[0]$ can be naturally identified with the Lie subalgebra of $\mathcal{L}$ from Section 1 via $j^{0}:D^{{\dagger}}_{qp^{M-1}a,n}\mapsto D_{an}$, where $a$ runs over $\mathbb{Z}^{0}(p)\cap[0,(p-1)v_{0})$. Then the special choice of $\delta_{0}$ and condition C3 from Section 2.1 imply that this identification is compatible with filtrations, i.e. for any $w\geqslant 1$, $j^{0}(\mathcal{L}^{{\dagger}}(w))=\mathcal{L}(w)$.
Let ¯$\bar{\mathcal{L}}^{\,{\dagger}}=\mathcal{L}^{\,{\dagger}}/\mathcal{L}^{{\dagger}}(p)$. This algebra is provided with the induced central filtration {¯$\{\bar{\mathcal{L}}^{\,{\dagger}}(w)\}_{w\geqslant 1}$ such that ¯$\bar{\mathcal{L}}^{\,{\dagger}}(p)=0$. We also have the induced gradings ¯$\bar{\mathcal{L}}^{\,{\dagger}}=\oplus_{m\geqslant 0}\bar{\mathcal{L}}^{\,{\dagger}}[m]$ and ¯$\bar{\mathcal{L}}^{\,{\dagger}}(w)=\oplus_{m\geqslant 0}\bar{\mathcal{L}}^{\,{\dagger}}(w)[m]$, where ¯$\bar{\mathcal{L}}^{\,{\dagger}}(w)[m]:=\bar{\mathcal{L}}^{\,{\dagger}}(w)\cap\bar{\mathcal{L}}^{\,{\dagger}}[m]$.
@@
If $l\in\bar{\mathcal{L}}^{\,{\dagger}}[m]_{W_{M}(k)}$, $l\neq 0$, we set $\mathrm{ch}(l)=m$.
@@
Clearly, for any $m_{1},m_{2}$, [¯$[\bar{\mathcal{L}}^{\,{\dagger}}[m_{1}],\bar{\mathcal{L}}^{\,{\dagger}}[m_{2}]]\subset\bar{\mathcal{L}}^{\,{\dagger}}[m_{1}+m_{2}]$ and we have the induced identification ¯$\bar{\mathcal{L}}^{{\dagger}}[0]=\bar{\mathcal{L}}$ which is compatible with filtrations.
2.4. Lie algebras $\mathcal{N}^{{\dagger}}$ and $\bar{\mathcal{N}}^{{\dagger}}$
Consider $N^{*},r^{*},q$ from Sec. 2.1.
@@ 2.4. Lie algebras $\mathcal{N}^{{\dagger}}$ and $\bar{\mathcal{N}}^{{\dagger}}$
Our approach to the study of the ramification filtration in $\mathcal{G}_{<p,M}$ substantially uses the construction of a totally ramified separable extension $\mathcal{K}^{\prime}$ of $\mathcal{K}$ such that $[\mathcal{K}^{\prime}:\mathcal{K}]=q$ and the Herbrand function $\varphi_{\mathcal{K}^{\prime}/\mathcal{K}}$ has only one edge point $(r^{*},r^{*})$. For a detailed construction of $\mathcal{K}^{\prime}$ cf. e.g. [3], Sect. 1.5. We just need that $\mathcal{K}^{\prime}=k((t_{01}))$, where our fixed local parameter $t_{0}$ of $\mathcal{K}$ equals $t_{01}^{q}E(t_{01}^{b^{*}})^{-1}$ and $E=E(X)$ is the Hasse-Artin exponential.
We use the uniformiser $t_{01}$ to construct (similarly to Sec. 1.1) the corresponding lifts $O^{\prime}_{M}(\mathcal{E}^{\prime})$ of separable extensions $\mathcal{E}^{\prime}/\mathcal{K}^{\prime}$. Notice that for any separable extension $\mathcal{E}^{\prime}$ of $\mathcal{K}^{\prime}$, $O^{\prime}_{M}(\mathcal{E}^{\prime})=W_{M}(\sigma^{M-1}\mathcal{E}^{\prime})[t_{1}]\subset W_{M}(\mathcal{E}^{\prime})$, where $t_{1}=[t_{01}]$ is the Teichmuller representative of $t_{01}$. In particular, $O^{\prime}_{M}(\mathcal{K}^{\prime})=W_{M}(k)((t_{1}))$. There is no natural embedding of $O_{M}(\mathcal{K})$ into $O^{\prime}_{M}(\mathcal{K}^{\prime})$. However, we have the embedding
$\iota:\sigma^{M-1}O_{M}(\mathcal{K})\subset O^{\prime}_{M}(\mathcal{K}^{\prime})$
@@
Let ${\mathcal{N}}^{\,{\dagger}}$ be the $W_{M}(k)$-submodule in ${\mathcal{L}}^{\,{\dagger}}_{O^{\prime}_{M}(\mathcal{K}^{\prime})}$ generated by the elements $p^{A}t_{1}^{-b}[D^{\,{\dagger}}_{\bar{\iota}\bar{n}}]$, where
a) $b\in\mathfrak{A}^{0}$, $\mathrm{ch}\,b\geqslant\mathrm{ch}(\bar{\iota})$;
b) $-b\geqslant-|\bar{\iota}|+p^{M-1}b^{*}U(\mathrm{ch}(\bar{\iota}),A)\,\,.$
Clearly, $\mathcal{N}^{{\dagger}}$ is a Lie $W_{M}(k)$-subalgebra in $\mathcal{L}^{{\dagger}}_{O^{\prime}_{M}(\mathcal{K}^{\prime})}$. We can define similarly $\bar{\mathcal{N}}^{{\dagger}}$ by replacing $\mathcal{L}^{{\dagger}}$ with $\bar{\mathcal{L}}^{{\dagger}}$. The natural projection $\mathcal{L}^{{\dagger}}\to\bar{\mathcal{L}}^{{\dagger}}$ induces the projection $\mathcal{N}^{{\dagger}}\to\bar{\mathcal{N}}^{{\dagger}}$.
@@
For $j,m\geqslant 0$, let ${\mathcal{N}}^{\,{\dagger}}[j,m]$ be the $W_{M}(k)$-submodule in ${\mathcal{N}}^{\,{\dagger}}$ generated by the elements $p^{A}t_{1}^{-b}[D^{\,{\dagger}}_{\bar{\iota}\bar{n}}]$, where
a) $\mathrm{ch}(\bar{\iota})=m$, $\mathrm{ch}\,b=j+m$;
b) $-b\geqslant-|\bar{\iota}|+p^{M-1}b^{*}U(m+j,A)\,.$
Clearly, there is a decomposition $\oplus_{j,m}{\mathcal{N}}^{\,{\dagger}}[j,m]={\mathcal{N}}^{\,{\dagger}}$ and
$[{\mathcal{N}}^{\,{\dagger}}[j_{1},m_{1}],{\mathcal{N}}^{\,{\dagger}}[j_{2},m_{2}]]\subset{\mathcal{N}}^{\,{\dagger}}[j_{1}+j_{2},m_{1}+m_{2}]\,.$

### v5

@@ 2.3. Lie algebras ${\mathcal{L}}^{\,{\dagger}}$ and $\bar{\mathcal{L}}^{\,{\dagger}}$
Let $\iota=p^{m}(qp^{M-1}\alpha-(q-1)\beta)\in\mathfrak{A}^{0}$ be given in notation from Sect. 2.2.
@@
The elements of $\{t^{-\iota}\ |\ \iota\in\mathfrak{A}^{+}(p)\}$ behave “well” modulo $(\sigma-\mathrm{id})\mathcal{K}$, i.e. the natural map $\sum\limits_{\iota\in\mathfrak{A}^{+}(p)}kt_{0}^{-\iota}\longrightarrow\mathcal{K}/(\sigma-\mathrm{id})\mathcal{K}$ is injective. This is implied by the following proposition.
@@
Indeed, $\iota p^{-m}=qp^{M-1}\alpha-(q-1)\beta\in p^{M}\mathbb{N}$ implies (use that $qp^{M-1}\alpha\in p^{M}\mathbb{N}$ because $m<N^{*}$) that $\beta/r^{*}=p^{-m}(b_{1}p^{m_{1}}+b_{2}p^{m_{2}}+\dots+b_{u}p^{m_{u}})\in p^{M}\mathbb{N}$ with all $m_{i}\leqslant m+M-1$. But this number is $\leqslant p^{M-1}\mathrm{ch}\,\iota<p^{M}$. The contradiction. Then Prop. 2.1a) implies that
@@
Finally, if $\iota\in\mathfrak{A}^{0}_{0}$ then $\iota p^{-v_{p}(\iota)}=a<(p-1)v_{0}$ implies the inequality $a<(v_{0}-\delta_{0})(p-1)$ by the choice of $\delta_{0}$, cf. Sect. 2.1. On the other hand, for all $\iota\in\mathfrak{A}^{+}(p)$ with $\mathrm{ch}(\iota)\geqslant 1$, the values $\iota p^{-v_{p}(\iota)}$ are different (use that $\mathrm{gcd}(p^{m}\alpha,p^{m}\beta/r^{*})\not\equiv 0\,\mathrm{mod}\,p$) and are bigger than $(v_{0}-\delta_{0})(p-1)$.
@@
Let ^$\widehat{\mathcal{L}}_{W(k)}^{\,{\dagger}}$ be the Lie $W(k)$-algebra with the set of free generators
@@
Use the notation $\sigma$ for the $\sigma$-linear automorphism of ^$\widehat{\mathcal{L}}^{\,{\dagger}}_{W(k)}$ such that $\sigma:D^{\,{\dagger}}_{\iota n}\mapsto D^{\,{\dagger}}_{\iota,n+1}$ and $\sigma:D^{{\dagger}}_{0}\mapsto D^{{\dagger}}_{0}$. Introduce the Lie $\mathbb{Z}_{p}$-algebras ^:=^$\widehat{\mathcal{L}}^{\,{\dagger}}:=\widehat{\mathcal{L}}^{\,{\dagger}}_{W(k)}|_{\sigma=\mathrm{id}}$ and =^$\mathcal{L}^{\,{\dagger}}=\widehat{\mathcal{L}}^{\,{\dagger}}/C_{p}(\widehat{\mathcal{L}}^{\,{\dagger}})\otimes\mathbb{Z}/p^{M}$.
@@
$\{\mathcal{L}^{\,{\dagger}}(w)\}_{w\geqslant 1}$ is the minimal filtration of $\mathcal{L}^{\,{\dagger}}$ such that:
– for all $w_{1},w_{2}$, $[\mathcal{L}^{{\dagger}}(w_{1}),\mathcal{L}^{{\dagger}}(w_{2})]\subset\mathcal{L}^{{\dagger}}(w_{1}+w_{2})$;
$p^{m}D^{\,{\dagger}}_{\iota n}\in\mathcal{L}^{\,{\dagger}}(1+s+\mathrm{ch}(\iota))_{W(k)}$ if $\iota\geqslant p^{M-1}b^{*}U(s,m)$.
@@
The above condition $\iota\geqslant b^{*}p^{M-1}U(s,m)$ is analogous to the condition $a\geqslant v_{0}U(s,m)$ from Section 1.4, where $v_{0}$ is replaced by $b^{*}/q$ (which is very close to $v_{0}$, cf. condition C3 from Section 2.1.
@@
For $\bar{\iota}=(\iota_{1},\dots,\iota_{r})$, set $r(\bar{\iota})=r$, $\mathrm{ch}(\bar{\iota})=\mathrm{ch}(\iota_{1})+\dots+\mathrm{ch}(\iota_{r})$ and $|\bar{\iota}|=\iota_{1}+\dots+\iota_{r}$.
Proceeding similarly to Sections 1.4-1.5 introduce the submodules $\mathcal{L}^{{\dagger}}(w)_{\mathcal{N}}$ in $\mathcal{L}^{{\dagger}}$ generated by $p^{m}[D^{{\dagger}}_{\bar{\iota}\bar{n}}]$ (this is an analogue of the notation $[D_{\bar{a}\bar{n}}]$) such that either $\mathrm{ch}(\bar{\iota})+1\geqslant w$ or $|\bar{\iota}|\geqslant p^{M-1}b^{*}U(w-\mathrm{ch}(\bar{\iota})-1,m)$, otherwise. Then we can prove an analogue of Corollary 1.7 that for any $1\leqslant w\leqslant p$,
(2.1) $\mathcal{L}^{{\dagger}}_{\mathcal{N}}(w)\subset\mathcal{L}^{{\dagger}}(w)\,.$
There is also an analogue of Proposition 1.4, where the quotients $a_{i}/v_{0}$ are replaced by $\iota_{i}/p^{M-1}b^{*}$ or, equivalently, by $\iota_{i}/qp^{M-1}v_{0}$.
More precisely, set
$M(\bar{\iota},s)=\min\{m\in\mathbb{Z}_{\geqslant 0}\ |\ p^{m}[D_{\bar{\iota}\bar{n}}]\in\mathcal{L}(r(\bar{\iota})+\mathrm{ch}(\bar{\iota})+s)_{W(k)}\}\,.$
We always assume that $r^{*}=r(\bar{\iota})+\mathrm{ch}(\bar{\iota})<p$ because $D_{\bar{\iota}\bar{n}}\in\mathcal{L}^{{\dagger}}(r^{*})_{W(k)}$.
If $\bar{a}=\bar{0}$ then, clearly, $M(\bar{a},0)=0$.
Assume that $\bar{a}\neq\bar{0}$ and $s\geqslant 1$.
Let $\bar{u}=\bar{u}(\bar{\iota})$ be the minimal integer such that for all $1\leqslant i\leqslant r$, there are $p$-digits $\theta_{i},\theta_{i}^{\prime}\in[0,p-1]$ such that
$\theta_{i}/p^{\bar{u}}+\theta_{i}^{\prime}/p^{\bar{u}+1}\leqslant\iota_{i}/qb^{*}<\theta_{i}/p^{\bar{u}}+(\theta_{i}^{\prime}+1)/p^{\bar{u}+1}\,,$
Since $\bar{u}$ is minimal, at least one $\theta_{i}\neq 0$.
Set $\bar{\kappa}_{0}=\kappa_{0}(\bar{\iota})=\sum_{1\leqslant i\leqslant r}\theta_{i}$ and $\bar{\kappa}_{1}=\kappa_{1}(\bar{\iota})=\sum_{1\leqslant i\leqslant r}\theta^{\prime}_{i}$.
Proposition 2.5.
a) $M(\bar{\iota},j)=j\bar{u}$ for $0\leqslant j\leqslant\min\{p-r^{*},\bar{\kappa}_{0}\}$;
b) $M(\bar{\iota},\bar{\kappa}_{0}+j)=\bar{u}\bar{\kappa}_{0}+(\bar{u}+1)j$, for $1\leqslant j\leqslant\min\{p-(r^{*}+\bar{\kappa}_{0}),\bar{\kappa}_{1}\}$;
c) $M(\bar{\iota},\bar{\kappa}_{0}+\bar{\kappa}_{1}+j)=\bar{\kappa}_{0}\bar{u}+\bar{\kappa}_{1}(\bar{u}+1)+1+j(\bar{u}+1)$, where $j$ runs over $1\leqslant j\leqslant p-(r^{*}+\bar{\kappa}_{0}+\bar{\kappa}_{1})$.
For $m\in\mathbb{Z}_{\geqslant 0}$, let $\mathcal{L}^{\,{\dagger}}[m]_{W(k)}$ be the $W(k)$-submodule of $\mathcal{L}^{{\dagger}}_{W(k)}$ generated by all monomials $[D^{\,{\dagger}}_{\bar{\iota}\bar{n}}]$ such that $\mathrm{ch}(\iota_{1})+\dots+\mathrm{ch}(\iota_{r})=m$. Denote by $\mathcal{L}^{{\dagger}}[m]$ the $\mathbb{Z}_{p}$-submodule of $\mathcal{L}^{{\dagger}}$ such that $\mathcal{L}^{{\dagger}}[m]\otimes W(k)=\mathcal{L}^{{\dagger}}[m]_{W(k)}$.
We obtained the grading $\mathcal{L}^{\,{\dagger}}=\oplus_{m\geqslant 0}\mathcal{L}^{\,{\dagger}}[m]$ in the category of Lie algebras. If $\mathcal{L}^{\,{\dagger}}(w)[m]=\mathcal{L}^{\,{\dagger}}(w)\cap\mathcal{L}^{\,{\dagger}}[m]$, then $\mathcal{L}^{\,{\dagger}}(w)=\oplus_{0\leqslant m\leqslant w-1}\mathcal{L}^{\,{\dagger}}(w)[m]$.
Remark.
The algebra $\mathcal{L}^{{\dagger}}[0]$ can be naturally identified with a Lie subalgebra of $\mathcal{L}$ from Section 1 via $j^{0}:D^{{\dagger}}_{qp^{M-1}a,n}\mapsto D_{an}$, where $a$ runs over the set $\mathbb{Z}^{0}(p)\cap[0,(p-1)v_{0})$. Then the special choice of $\delta_{0}$ and condition C3 from Section 2.1 imply that this identification is compatible with filtrations, i.e. for any $w\geqslant 1$, $j^{0}(\mathcal{L}^{{\dagger}}(w)[0])\subset\mathcal{L}(w)$.
Let ¯$\bar{\mathcal{L}}^{\,{\dagger}}=\mathcal{L}^{\,{\dagger}}/\mathcal{L}^{{\dagger}}(p)$. This algebra is provided with the induced central filtration {¯$\{\bar{\mathcal{L}}^{\,{\dagger}}(w)\}_{w\geqslant 1}$ such that ¯$\bar{\mathcal{L}}^{\,{\dagger}}(p)=0$. We also have the induced gradings ¯$\bar{\mathcal{L}}^{\,{\dagger}}=\oplus_{0\leqslant m\leqslant p-2}\bar{\mathcal{L}}^{\,{\dagger}}[m]$ and ¯$\bar{\mathcal{L}}^{\,{\dagger}}(w)=\oplus_{0\leqslant m\leqslant w-1}\bar{\mathcal{L}}^{\,{\dagger}}(w)[m]$, where ¯$\bar{\mathcal{L}}^{\,{\dagger}}(w)[m]:=\bar{\mathcal{L}}^{\,{\dagger}}(w)\cap\bar{\mathcal{L}}^{\,{\dagger}}[m]$. Note that ¯$\bar{\mathcal{L}}^{{\dagger}}[p-1]=0$.
@@
Clearly, for any $m_{1},m_{2}$, [¯$[\bar{\mathcal{L}}^{\,{\dagger}}[m_{1}],\bar{\mathcal{L}}^{\,{\dagger}}[m_{2}]]\subset\bar{\mathcal{L}}^{\,{\dagger}}[m_{1}+m_{2}]$ and we have the induced identification ¯$\bar{\mathcal{L}}^{{\dagger}}[0]=\bar{\mathcal{L}}$ which is compatible with filtrations. (Use that all $D_{an}$ with $a\geqslant(p-1)v_{0}$ belong to ${\mathcal{L}}(p)_{W(k)}$.)
2.4. Lie algebras $\mathcal{N}^{{\dagger}}$, $\bar{\mathcal{N}}^{{\dagger}}$ and $\widetilde{\mathcal{N}}^{{\dagger}}$
Consider $N^{*},r^{*},q$ from Section 2.1.
@@
Our approach to the study of the ramification filtration in $\mathcal{G}_{<p,M}$ substantially uses the construction of a totally ramified separable extension $\mathcal{K}^{\prime}$ of $\mathcal{K}$ such that $[\mathcal{K}^{\prime}:\mathcal{K}]=q$ and the Herbrand function $\varphi_{\mathcal{K}^{\prime}/\mathcal{K}}$ has only one edge point $(r^{*},r^{*})$. For a detailed construction of $\mathcal{K}^{\prime}$ cf. e.g. [3], Sect. 1.5. We just need that $\mathcal{K}^{\prime}=k((t_{01}))$, where our fixed local parameter $t_{0}$ of $\mathcal{K}$ equals $t_{01}^{q}E(t_{01}^{b^{*}})^{-1}$ and $E=E(X)$ is the Artin-Hasse exponential.
We do not need here a general concept of a lift modulo $p^{M}$ which was used e.g. in [2]. In the case of the field $\mathcal{K}$ its lift $O_{M}(\mathcal{K})$ with respect to the parameter $t_{0}$ appears as $W_{M}(k)((t))\subset W_{M}(\mathcal{K})$, where $t=[t_{0}]$ is the Teichmuller representative of $t_{0}$. We need also a similar lift for $\mathcal{K}^{\prime}$ related to its local parameter $t_{01}$. So, we set $O^{\prime}_{M}(\mathcal{K}^{\prime})=W_{M}(k)((t_{1}))\subset W_{M}(\mathcal{K}^{\prime})$, where $t_{1}=[t_{01}]$. There is no a natural embedding of $O_{M}(\mathcal{K})$ into $O^{\prime}_{M}(\mathcal{K}^{\prime})$. However, we have the embedding
$\iota:\sigma^{M-1}O_{M}(\mathcal{K})=W_{M}(k)((t^{p^{M-1}}))\subset O^{\prime}_{M}(\mathcal{K}^{\prime})$
@@
Let ${\mathcal{N}}^{\,{\dagger}}$ be the $W(k)$-submodule in ${\mathcal{L}}^{\,{\dagger}}_{O^{\prime}_{M}(\mathcal{K}^{\prime})}$ generated by the elements $t_{1}^{-b}[D^{\,{\dagger}}_{\bar{\iota}\bar{n}}]$, where $-b\geqslant-|\bar{\iota}|$.
Clearly, $\mathcal{N}^{{\dagger}}$ is a Lie $W(k)$-subalgebra in $\mathcal{L}^{{\dagger}}_{O_{M}^{\prime}(\mathcal{K}^{\prime})}$. We can define similarly $\bar{\mathcal{N}}^{{\dagger}}$ by replacing $\mathcal{L}^{{\dagger}}$ to $\bar{\mathcal{L}}^{{\dagger}}$. Then a natural projection $\mathcal{L}^{{\dagger}}\to\bar{\mathcal{L}}^{{\dagger}}$ induces the projection $\mathcal{N}^{{\dagger}}\to\bar{\mathcal{N}}^{{\dagger}}$.
@@
For $s\geqslant 0$,
a) let ${\mathcal{N}}^{\,{\dagger}}\langle s\rangle$ be the $W(k)$-submodule in ${\mathcal{N}}^{\,{\dagger}}$ generated by all $p^{A}t_{1}^{-b}[D^{\,{\dagger}}_{\bar{\iota}\bar{n}}]$, where $-b\geqslant-|\bar{\iota}|+p^{M-1}b^{*}U(s,A)\,;$
b) let ${\mathcal{N}}^{\,{\dagger}}(s)$ be the $W(k)$-submodule in ${\mathcal{N}}^{\,{\dagger}}$ generated by all ${\mathcal{N}}^{\,{\dagger}}\langle s_{1}\rangle[m]$ with $s_{1}+m=s$.
We obtained a decreasing filtration $\{\mathcal{N}^{\,{\dagger}}(s)\}_{s\geqslant 0}$ of $\mathcal{N}^{{\dagger}}$ such that
$[{\mathcal{N}}^{\,{\dagger}}(s_{1}),{\mathcal{N}}^{\,{\dagger}}(s_{2})]\subset{\mathcal{N}}^{\,{\dagger}}(s_{1}+s_{2})\,.$
Proposition 2.6.
For any $s,m$ and $\mathrm{m}^{\prime}=t_{1}W_{M}(k)[[t_{1}]]$, it holds
$\mathcal{N}^{{\dagger}}\langle s\rangle[m]\subset\mathcal{L}^{{\dagger}}_{\mathrm{m}^{\prime}}+\mathcal{L}^{{\dagger}}(1+m+s)_{O_{M}^{\prime}(\mathcal{K}^{\prime})}\,.$
Proof.
If $p^{A}t_{1}^{-b}[D^{{\dagger}}_{\bar{\iota}\bar{n}}]\in\mathcal{N}^{{\dagger}}\langle s\rangle[m]$ then $-b\geqslant-|\bar{\iota}|+p^{M-1}b^{*}U(s,A)$. If $|\bar{\iota}|\geqslant p^{M-1}b^{*}U(s,A)$ then $p^{A}[D^{{\dagger}}_{\bar{\iota}\bar{n}}]\in\mathcal{L}^{{\dagger}}(1+m+s)_{W(k)}\,.$ Otherwise, $-b\geqslant 1$. ∎
Similar definitions can be done in the context of the Lie algebra $\bar{\mathcal{L}}^{{\dagger}}$ and the analog $\bar{\mathcal{N}}^{{\dagger}}$ of $\mathcal{N}^{{\dagger}}$ with the corresponding filtration $\bar{\mathcal{N}}^{{\dagger}}(s)$. The projection $\mathcal{N}^{{\dagger}}\to\bar{\mathcal{N}}^{{\dagger}}$ allows us to transform above properties to $\bar{\mathcal{N}}^{{\dagger}}$. In particular, we have the following
Corollary 2.7.
a) $\bar{\mathcal{N}}^{{\dagger}}(p-1)\subset\bar{\mathcal{L}}^{{\dagger}}_{\mathrm{m}^{\prime}}$;
b) if $p^{A}t^{-b}[D^{\,{\dagger}}_{\bar{\iota}\bar{n}}]\in\bar{\mathcal{N}}^{\,{\dagger}}$ is non-zero then $|\bar{\iota}|\in\mathfrak{A}^{0}$;
c) if $p^{A}t^{-b}[D^{\,{\dagger}}_{\bar{\iota}\bar{n}}]\notin\bar{\mathcal{N}}^{\,{\dagger}}(p-1)$ then $-b\leqslant p^{M-1}(p-1)b^{*}$.